# Proximal Policy Distillation

### v1

@@ 6 Acknowledgments
We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius. This project was supported by SURF grant EINF-5635.
@@ B.2 Policy distillation
We then report individual scores for all environments and students in Table 5.
@@ B.3 Distillation: imperfect teachers
Table 6: Full results for each environment, extending Table 2 from the main text. We show the performance of student models, trained using the three distillation methods (PPD, student-distill, and teacher-distill) from ‘imperfect teachers’ that are artificially corrupted to decrease in performance. Results are calculated as a fraction of the original teacher score for each environment and random seed, and then averaged by geometric mean. Distillation is performed on four Atari and four Procgen environments, and onto larger student networks. Evaluation for Procgen is performed on test levels. We omit the ’x’ symbols to reduce clutter.
@@
BeamRider 3754.15 0.29x 0.33 0.31 0.39
CrazyClimber 62912.0 0.39x 0.37 0.44 0.54
MsPacman 2640.16 0.49x 0.52 0.55 0.71
Qbert 7904.5 0.54x 0.66 0.6 0.81
atari 0.41x 0.45x 0.46x 0.59x
miner 6.64 0.69x 0.36 0.48 0.69
jumper 8.39 0.56x 0.6 0.63 0.63
coinrun 7.29 0.8x 0.71 0.77 0.83
ninja 6.38 0.64x 0.69 0.65 0.72
procgen 0.67x 0.57x 0.63x 0.71x
B.4 Procgen: train vs test levels
@@
Table 7: Performance of students models of three sizes (smaller, same, larger), trained using the three distillation methods (PPD, student-distill, and teacher-distill). Evaluation is performed separately on the train and test levels of Procgen, contrary to Table 1 that only shows performance evaluated on unseen test levels. Results are calculated as a fraction of teacher score for each environment and random seed, and then averaged by geometric mean.
env teacher smaller same-size larger
score td sd ppd td sd ppd td sd ppd
procgen (train) 7.17 0.5x 0.55x 0.64x 0.67x 0.71x 0.85x 0.76x 0.83x 0.93x
procgen (test) 5.16 0.59x 0.66x 0.75x 0.75x 0.82x 0.98x 0.83x 0.93x 1.02x
Table 8: Performance of student models, trained using the three distillation methods (PPD, student-distill, and teacher-distill) from ‘imperfect teachers’ that are artificially corrupted to decrease in performance. Results are calculated as a fraction of the original teacher score for each environment and random seed, and then averaged by geometric mean. Distillation is performed on four Atari and four Procgen environments, and onto larger student networks. We omit the ’x’ symbols to reduce clutter. Evaluation is performed on training levels.
env original teacher corrupted teacher larger
@@
miner 6.64 0.82x 0.41 0.56 0.82
jumper 8.39 0.89x 0.87 0.88 0.91
coinrun 7.29 0.88x 0.86 0.88 0.92
ninja 6.38 0.78x 0.78 0.75 0.83
procgen 0.84x 0.7x 0.76x 0.87x

### v2

@@ 6 Acknowledgments
This research was supported by SURF grant EINF-5635. We gratefully acknowledge SURF (https://www.surf.nl) for providing access to the National Supercomputer Snellius. We also thank Murat Kirtay and Bosong Ding for their valuable support and stimulating discussions, and the anonymous reviewers for their insightful feedback, which substantially improved this manuscript.
@@ A.2 Training of the teacher models
total env steps 10 million (Atari and Mujoco), 30 million (Procgen)
@@ B.2 Policy distillation
We then report individual scores for all environments and students in Table 5, and the individual, per-environment training trajectories for distillation to larger student networks for PPD and student-distill in Figures 9, 10, and 11, that is, the training curves that are combined to form Figure. 1 in the main text..
@@
Figure 9: Individual per-environment (Atari) training curves during distillation with PPD and student-distill onto larger student networks. Figure 1 (left) combines this data obtained by interquartile averaging. Curves shown are the interquartile mean over 5 random seeds. Shaded areas denote 95% confidence intervals.
Figure 10: Individual per-environment (Mujoco) training curves during distillation with PPD and student-distill onto larger student networks. Figure 1 (left) combines this data obtained by interquartile averaging. Curves shown are the interquartile mean over 5 random seeds. Shaded areas denote 95% confidence intervals.
Figure 11: Individual per-environment (Procgen) training curves during distillation with PPD and student-distill onto larger student networks. Figure 1 (left) combines this data obtained by interquartile averaging. Curves shown are the interquartile mean over 5 random seeds. Shaded areas denote 95% confidence intervals.
@@ B.3 Distillation: imperfect teachers
Table 6: Full results for each environment, extending Table 2 from the main text. We show the performance of student models, trained using the three distillation methods (PPD, student-distill, and teacher-distill) from ‘imperfect teachers’ that are artificially corrupted to decrease in performance. Results are calculated as a fraction of the original teacher score for each environment and random seed, and then aggregated by geometric interquartile mean (95% confidence intervals in brackets). Distillation is performed on four Atari (human-normalized scores) and four Procgen environments, and onto larger student networks.
@@
BeamRider 0.2 0.15x 0.14x 0.16x 0.19x
CrazyClimber 2.08 0.22x 0.22x 0.3x 0.41x
MsPacman 0.35 0.38x 0.43x 0.47x 0.66x
Qbert 0.58 0.52x 0.65x 0.59x 0.8x
atari 0.42x [0.33, 0.57] 0.48x [0.38, 0.61] 0.52x [0.38, 0.68] 0.69x [0.55, 0.82]
miner 6.64 0.69x 0.36x 0.48x 0.69x
jumper 8.39 0.56x 0.6x 0.63 0.63x
coinrun 7.29 0.8x 0.71x 0.77x 0.83x
ninja 6.38 0.64x 0.69x 0.65x 0.72x
procgen 0.65x [0.63, 0.69] 0.63x [0.59, 0.67] 0.64x [0.61, 0.67] 0.71x [0.68, 0.74]
B.4 PPD versus PPO
We analyze the effectiveness of combining distillation and reinforcement learning in PPD by comparing against training student policies with pure PPO (without the distillation loss). We perform this comparison using larger student networks on the same subset of environments used in the imperfect teachers experiments (four Procgen and four Atari environments).
The results, shown in Table 7, demonstrate that PPD substantially outperforms pure PPO training, highlighting the value of incorporating teacher guidance through our proximal distillation approach.
Table 7: Baseline comparison of PPO (without distillation) and the three distillation methods compared in our work: teacher-distill, student-distill, and PPD. Results are calculated as fraction of teacher score for each environment and random seed, and then aggregated by geometric interquartile mean (95% confidence intervals are included in brackets).
env teacher larger
score ppo td sd ppd
atari 0.48 0.38x [0.27, 0.51] 1.13x [1.02, 1.28] 1.13x [1.03, 1.29] 1.16x [1.05, 1.33]
procgen (train) 7.07 0.44x [0.38, 0.5] 0.86x [0.83, 0.89] 0.86x [0.83, 0.9] 0.95x [0.92, 0.98]
procgen (test) 5.07 0.71x [0.63, 0.79] 0.91x [0.86, 0.98] 0.96x [0.93, 1.0] 1.04x [0.97, 1.08]
B.5 Procgen: train vs test levels
@@
Table 8: Performance of students models of three sizes (smaller, same, larger), trained using the three distillation methods (PPD, student-distill, and teacher-distill). Evaluation is performed separately on the train and test levels of Procgen, contrary to Table 1 that only shows performance evaluated on unseen test levels. Results are calculated as a fraction of teacher score for each environment and random seed, and then aggregated by geometric interquartile mean (95% confidence intervals in brackets).
teacher smaller
score td sd ppd
procgen (train) 7.07 0.62x [0.6, 0.66] 0.71x [0.66, 0.76] 0.81x [0.77, 0.87]
procgen (test) 5.07 0.72x [0.69, 0.77] 0.78x [0.74, 0.83] 0.93x [0.87, 1.01]
teacher same-size
score td sd ppd
procgen (train) 7.07 0.76x [0.74, 0.79] 0.79x [0.75, 0.83] 0.88x [0.84, 0.91]
procgen (test) 5.07 0.84x [0.78, 0.89] 0.88x [0.84, 0.94] 1.03x [0.97, 1.08]
teacher larger
score td sd ppd
procgen (train) 7.07 0.86x [0.83, 0.89] 0.86x [0.83, 0.9] 0.95x [0.92, 0.98]
procgen (test) 5.07 0.91x [0.86, 0.98] 0.96x [0.93, 1.0] 1.04x [0.98, 1.08]
Table 9: Performance of student models, trained using the three distillation methods (PPD, student-distill, and teacher-distill) from ‘imperfect teachers’ that are artificially corrupted to decrease in performance. Results are calculated as a fraction of the original teacher score for each environment and random seed, and then aggregated by geometric interquartile mean (95% confidence intervals in brackets). Distillation is performed on four Atari and four Procgen environments, and onto larger student networks. Evaluation is performed on training levels.
env original corrupted larger
@@
miner 6.64 0.82x 0.41x 0.56x 0.82x
jumper 8.39 0.89x 0.87x 0.88x 0.91x
coinrun 7.29 0.88x 0.86x 0.88x 0.92x
ninja 6.38 0.78x 0.78x 0.75x 0.83x
procgen (train) 0.86x [0.79, 0.91] 0.79x [0.77, 0.83] 0.79x [0.75, 0.83] 0.87x [0.84, 0.89]